# Umberto Collodel

### v2

@@ 4.2 Few-Shot Prompt Results
The few-shot specification does not materially improve correlation with market volatility compared to the naive model: correlations remain essentially unchanged across maturities, 2-year (0.52 vs. 0.48), 10-year (0.45 vs. 0.46), and 3-month (0.38 vs. 0.38), and none of these differences are statistically significant. The directional responses likewise mirror those produced under the zero-shot specification. Where the few-shot treatment matters is, however, in magnitude calibration. As Figure 7 shows, providing shared historical examples compresses forecast dispersion sharply, with maximum uncertainty for longer tenors falling from roughly 0.6 to about 0.3 percentage points, a level comparable to observed market-based volatility (Figure 1). This improved calibration of cross-sectional variance also makes the model’s first-moment forecasts much closer to reality.
@@
Note: Each data point reflects the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025), prompt 2 and provides the text of the three previous conferences as well as the ensuing market-based standard deviation of OIS rates. Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
@@ 4.3 LLM-as-a-Judge
Note: Each point reflects the average Spearman correlation across tenors (3months, 2-years and 10-years) between LLM-generated rate Euro OIS forecasts and market-based monetary policy uncertainty from Collodel and Kunzmann, (2025) after ECB GovC press conferences.The x-axis indicates the number of prompt iteration in a LLM-as-Judge framework. The simulation uses Gemini 2.5-Flash with a temperature of 1, and incorporates a feedback loop via a separate LLM (“Judge”) that uses Gemini 2.5-Pro. Train and test set are 75/25.
@@
Note: Time series comparison of LLM-predicted forecast dispersion and actual market volatility following ECB press conferences, segmented by prompt iteration and tenor. Simulations employ Gemini 2.5-Flash with temperature = 1 and a Judge-based feedback loop that iteratively refines prompts over 5 iterations. Forecast dispersion represents the standard deviation across 5 simulation runs per prompt per conference. Market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025). Red shaded areas indicate periods where prediction errors exceed the 90th percentile. Spearman correlation coefficients (r) between predicted and actual volatility are shown in red for each panel.
@@ 5 Robustness
To validate our findings, we evaluate the robustness of the baseline LLM-based disagreement measure across three complementary dimensions that address key methodological concerns. First, we compare our approach with simple textual benchmarks to establish whether sophisticated agent-based modeling offers genuine advantages over basic complexity metrics. Second, we assess model architecture stability across different LLM systems (Gemini, ChatGPT, and Claude) to ensure our results reflect systematic patterns captured by this class of models rather than artifacts of a particular model’s design. Third, we examine prompt specification stability, addressing the well-documented fickleness of LLMs to minor wording changes.
@@ 5.1 Simple Benchmark Comparison
A critical question to justify our approach is whether sophisticated LLM-based agent modeling offers genuine advantages over simple textual measures, or whether basic complexity metrics can achieve similar predictive power at much lower computational cost. To address this fundamental methodological question, we compare the LLM-based correlations with a simple textual benchmark: the Flesch-Kincaid readability score. The Flesch-Kincaid score is a standard measure of text readability that estimates the U.S. school grade level required to understand a given text, with higher scores indicating more complex language.
In related work, Mumtaz et al., (2023) show that communication complexity at the Bank of England is associated with higher gilt yield volatility, with complex statements producing significantly larger increases in the standard deviation of returns compared to simpler ones. Our analysis tests whether this relationship holds for ECB communications and how it compares to our LLM-based approach.
Figure 10: Correlation between Flesh-Kincaid score and post ECB conference OIS volatility
Note: Volatility in percentage points.
The correlations between Flesch-Kincaid scores and post-conference OIS volatility are 0.15, 0.26, and 0.06 for the 3-month, 2-year, and 10-year tenors, respectively (Figure 10). These results suggest that readability and complexity might be part of the story: more complex communication can amplify market uncertainty. Nevertheless, this dimension captures only one small facet of the broader interpretive process. By construction, readability indices cannot capture policy content, such as the magnitude of the monetary policy surprise, or strategy shifts, like the abandonment of forward guidance, so their explanatory power remains limited. For example, during the Covid-19 pandemic, policy statements became highly complex owing to the unprecedented nature of the health shock, yet forward guidance kept expectations well anchored and first-moment surprises remained small, thereby limiting volatility. In contrast, our LLM-based framework delivers substantially higher correlations by capturing the heterogeneous interpretive processes through which market participants react to central bank communication.
@@ 6 Conclusions
This paper introduces a novel methodology using Large Language Models to simulate heterogeneous market reactions to ECB press conferences. Through an agent-based framework with 30 synthetic traders, each endowed with distinct behavioral characteristics, we demonstrate that LLMs can effectively capture market disagreement patterns in response to monetary policy communication.
Our key findings advance both theoretical understanding and practical application in monetary economics. First, even a naive zero-shot approach achieves substantial correlations (roughly 0.5) with actual market disagreement, particularly for medium and long-term maturities. This suggests LLMs can effectively decode central bank communication in economically meaningful ways. Second, the LLM-as-a-Judge framework substantially enhances performance at early iterations (average correlation across tenors reaching 0.57), though subsequent iterations show degradation, emphasizing the need for human oversight in automated prompt optimization. The stability of results across training and test datasets demonstrates that our approach captures fundamental communication-market relationships rather than memorizing patterns.
These results have significant practical implications. For central banks, our framework offers a powerful ex-ante tool to evaluate communication strategies and identify potentially market-disrupting language before release. For market participants, it provides insights into the drivers of post-announcement volatility. For researchers, it opens the “black box” of expectation formation by providing a micro-founded framework that explicitly models interpretive processes, moving beyond traditional reduced-form event studies and rule-based models. The methodology is flexible and readily adaptable to other central banks and policy contexts.
Despite these contributions, some limitations remain. The closed-source nature of leading LLMs constrains replicability, as future model versions may yield slightly different outputs. While we mitigate this problem through multi-model robustness testing, future research could explore open-source alternatives for greater transparency and reproducibility. Regarding potential contamination, while ECB press conferences are relatively niche compared to mainstream internet content, the possibility that fragments of such communication, or analyses thereof, were included in the model’s pretraining data cannot be completely dismissed. We explicitly instruct the model to use only information available up to each conference date; however, this constraint affects responses, not underlying parameters. Nevertheless, we consider substantial contamination risk limited given that discussions of post-ECB press conference volatility are rare in the public domain. Finally, domain-specific or fine-tuned models could potentially enhance performance by balancing interpretive diversity with specialized financial expertise, though this remains to be tested.
In conclusion, this paper establishes a new paradigm for analyzing monetary policy communication at the intersection of artificial intelligence, behavioral finance, and central banking practice. By showing that LLMs can simulate heterogeneous trader interpretations and capture market disagreement patterns, we provide policymakers with a powerful tool to anticipate market reactions and refine communication strategies, while offering researchers a framework to better understand the micro-foundations of expectation formation.

### v3

@@ 4.2 Few-Shot Prompt Results
The few-shot specification does not materially improve correlation with market volatility compared to the naive model: correlations remain essentially unchanged across maturities, 2-year (0.53 vs. 0.48), 10-year (0.44 vs. 0.46), and 3-month (0.39 vs. 0.38), and none of these differences are statistically significant. The directional responses likewise mirror those produced under the zero-shot specification. Where the few-shot treatment matters is, however, in magnitude calibration. As Figure 7 shows, providing shared historical examples compresses forecast dispersion sharply, with maximum uncertainty for longer tenors falling from roughly 0.6 to about 0.3 percentage points, a level comparable to observed market-based volatility (Figure 1). This improved calibration of cross-sectional variance also makes the model’s first-moment forecasts much closer to reality.
@@
Note: Each data point reflects the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025), the baseline prompt (Figure 2), and provides the text of the three previous conferences as well as the ensuing market-based standard deviation of OIS rates. Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
@@ 4.3 LLM-as-a-Judge
Note: Each point reflects the average Spearman correlation across tenors (3months, 2-years and 10-years) between LLM-generated standard deviation of Euro OIS rate forecasts and market-based monetary policy uncertainty from Collodel and Kunzmann, (2025) after ECB GovC press conferences.The x-axis indicates the number of prompt iteration in a LLM-as-Judge framework. The simulation uses Gemini 2.5-Flash with a temperature of 1, and incorporates a feedback loop via a separate LLM (“Judge”) that uses Gemini 2.5-Pro. Train and test set are 75/25.
@@
Note: Time series comparison of LLM-predicted forecast dispersion and actual market volatility following ECB press conferences, segmented by prompt iteration and tenor. Simulations employ Gemini 2.5-Flash with temperature = 1 and the second prompt generated by a Judge-based feedback loop that iteratively refines prompts over 5 iterations. Forecast dispersion represents the standard deviation across 5 simulation runs per prompt per conference. Market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025). Red shaded areas indicate periods where prediction errors exceed the 90th percentile. Spearman correlation coefficients (r) between predicted and actual volatility are shown in red for each panel.
@@ 5 Robustness
To validate our findings, we evaluate the robustness of the baseline LLM-based disagreement measure across three complementary dimensions that address key methodological concerns. First, we compare our approach with simple textual benchmarks to establish whether sophisticated agent-based modeling offers genuine advantages over basic complexity metrics. Second, we assess model architecture stability across different LLM systems (Gemini, ChatGPT, and Claude) to ensure our results reflect systematic patterns captured by this class of models rather than artifacts of a particular model’s design. Third, we examine prompt specification stability, addressing the well-documented fickleness of LLMs to minor wording changes (Loya et al.,, 2023; Sclar et al.,, 2024).
@@ 5.1 Simple Benchmark Comparison
A critical question to justify our approach is whether sophisticated LLM-based agent modeling offers genuine advantages over simple textual measures, or whether basic complexity metrics can achieve similar predictive power at much lower computational cost. To address this fundamental methodological question, we compare the LLM-based correlations with a simple textual benchmark: the Flesch-Kincaid readability score. The Flesch-Kincaid score is a standard measure of text readability that estimates the U.S. school grade level required to understand a given text, with higher scores indicating more complex language and longer, more technical sentences. In related work, Mumtaz et al., (2023) show that communication complexity at the Bank of England is associated with higher gilt yield volatility, with complex statements producing significantly larger increases in the standard deviation of returns compared to simpler ones. Our analysis tests whether this relationship holds for ECB communications and how it compares to our LLM-based approach.
Figure 10: Flesch-Kincaid complexity and post-ECB conference OIS volatility
Note: Market volatility measured as the average min-max range of OIS rates in the 3 days following ECB press conferences. Outliers (greater than 99th percentile) excluded. LOESS smoothing reflects rank-based relationship. Spearman correlation ($\rho$) shown in the top-right corner of each panel.
Figure 10 presents the relationship between Flesch-Kincaid complexity scores and market-based disagreement across maturities. The Spearman correlations are modest: 0.23 for both 3-month and 2-year tenors, and essentially zero (0.03) for the 10-year tenor. These results suggest that while syntactic complexity may play a role, it explains only a small fraction of market disagreement. Moreover, the relationship appears non-linear, particularly at the 2-year maturity, where disagreement peaks at moderate complexity levels before declining, a pattern inconsistent with a simple ”more complexity equals more uncertainty” narrative.
This weak correlation between readability metrics and market outcomes highlights a fundamental limitation: by construction, readability indices measure only surface-level linguistic features, such as sentence length, word difficulty, and grammatical structure, but cannot capture the substantive dimensions that drive interpretive disagreement i.e. policy content (the magnitude of monetary surprises), strategic shifts (such as the abandonment of forward guidance), internal contradictions between prepared statements and Q&A sessions, or the excessive use of hedging language. A paramount example is the COVID-19 pandemic: during this period, policy statements became highly complex due to the unprecedented nature of the health shock; yet, strong forward guidance kept expectations well anchored, and first-moment surprises remained small, thereby limiting volatility despite high Flesch-Kincaid scores.
In stark contrast, our LLM-based framework achieves correlations of approximately 0.5 for medium- and long-term tenors, more than double the readability-based benchmark. This substantial improvement reflects the LLM’s ability to capture what readability metrics cannot: how heterogeneous agents actually interpret policy signals, decode ambiguous forward guidance, and react to conflicting messages within the same communication. These findings validate our core methodological contribution: understanding market reactions to central bank communication requires modeling the interpretive processes of market participants, not merely measuring syntactic complexity.
@@ 6 Conclusions
Central banks face a fundamental communication challenge: they cannot test how markets will interpret their statements before release. Traditional high-frequency identification methods measure market reactions ex-post, revealing what happened but offering no guidance for ex-ante communication design. This constraint is particularly costly given mounting evidence that interpretive disagreement affects financial stability and monetary transmission. This paper addresses the gap by developing the first ex-ante framework for simulating heterogeneous market reactions to central bank communication using Large Language Models. We construct an agent-based simulation of 30 synthetic traders with distinct behavioral characteristics who process ECB press conference transcripts and forecast Euro interest rate swap rates across the yield curve. Cross-sectional forecast dispersion provides a model-based measure of market disagreement that can be computed before any communication event occurs.
Our empirical analysis of 283 ECB press conferences (June 1998–April 2025) yields three main findings. First, even without historical calibration, the zero-shot specification achieves Spearman correlations of approximately 0.5 with realized OIS volatility, demonstrating that LLMs possess intrinsic capabilities to decode central bank communication in economically meaningful ways. Second, the relationship proves strongest for the 2-year tenor ($\rho$ = 0.53), consistent with this maturity’s elevated sensitivity to forward guidance that dominated ECB communication for much of the period under review. The hierarchy across maturities, with medium-term rates most predictable, followed by long-term, then short-term, reflects fundamental differences in how policy signals transmit across the yield curve. Third, an adaptive “LLM-as-a-Judge” framework raises correlations to 0.6 at early iterations, though subsequent refinements degrade performance, demonstrating both the potential and limits of automated prompt optimization without human oversight.
These findings carry direct policy implications. For central banks, the framework provides an operational tool to anticipate communication-induced volatility before release, enabling systematic evaluation of alternative phrasings, identification of potentially ambiguous language, and quantitative assessment of how textual changes would affect expected market disagreement (refer to section Appendix C: Counterfactual Communication Analysis in Appendix C for a simple counterfactual analysis). This transforms the communication strategy from reactive refinement, based on costly market reactions, to proactive optimization during the drafting process. For market participants, the methodology offers insights into drivers of post-announcement volatility and a framework for evaluating their own interpretive biases. For researchers, it establishes a micro-founded approach to studying expectation formation that explicitly models heterogeneous belief processes, moving beyond reduced-form event studies. The methodology’s adaptability extends its relevance beyond the ECB. The framework transfers naturally to other major central banks, Federal Reserve, Bank of England, Bank of Japan, enabling comparative analysis of how communication styles, institutional frameworks, and forward guidance regimes affect interpretive disagreement across monetary policy contexts. The broader implications extend to how we conceptualize the role of AI in economic research and policy. Our results demonstrate that LLMs can serve not merely as statistical pattern recognizers but as computational representations of human interpretive processes, in line with Horton, (2023).
Nevertheless, important limitations warrant acknowledgment and suggest directions for future research. First, replicability faces constraints from LLM versioning and closed-source architectures, though our cross-model robustness checks and public release of prompts partially mitigate these concerns. Future work should prioritize open-source implementations for full transparency and reproducibility. Second, while we cannot completely rule out training data contamination, fragments of ECB communication or related analyses may have appeared in pretraining corpora, the rarity of detailed discussions about post-ECB volatility in public sources, combined with our temporal instructions limiting the model to information available at each conference date, substantially reduces this risk. Third, our reliance on general-purpose models may understate achievable performance. Domain-specific fine-tuning on central bank and financial text, or development of specialized architectures optimized for monetary policy interpretation, could significantly enhance accuracy.
In conclusion, this paper establishes that LLM-based agent simulations can quantitatively reproduce market disagreement patterns arising from central bank communication with economically significant accuracy. By demonstrating that interpretive heterogeneity is both measurable and predictable from text alone, we provide central banks with a practical tool for ex-ante communication evaluation and offer researchers a framework for studying the micro-foundations of expectation formation in monetary policy transmission.
@@ 15 Medium Position-taking behavior; different output headers Task & Output
Appendix C: Counterfactual Communication Analysis
6.5 Motivation and Design
Having established that LLM-based agents can provide central banks with an ex-ante assessment of the disagreement their communication is likely to generate, we now turn to a complementary question: Can central banks leverage LLMs to systematically refine communication drafts and thereby reduce ensuing interpretive uncertainty?
To answer this question, we design a simple counterfactual experiment. The experiment proceeds in three stages. First, we identify 30 high-volatility press conferences from our sample, defined as those with the highest average post-announcement OIS volatility. These high-volatility episodes are precisely the communications that matter the most for central banks, making them the ideal sample for evaluating whether linguistic clarity through LLM refinement can reduce market uncertainty. Second, we employ ChatGPT 5o-mini to edit the text of these conferences, with explicit instructions to reduce ambiguity while maintaining the substantive policy message (Figure 18). Third, we submit both the original and edited transcripts to our Gemini 2.5-Flash simulation framework, generating synthetic disagreement measures for each version. The change in disagreement between the two versions, measured both in percentage terms and absolute basis points, serves as our Average Treatment Effect (ATE).
Figure 18: ChatGPT 5o-mini Editing Prompt
Counterfactual Editing Prompt You are editing an ECB press conference to reduce ONLY linguistic ambiguity. STRICT CONSTRAINTS - DO NOT CHANGE: - Any policy decisions (interest rates, asset purchases, programs) - Any economic data, statistics, or forecasts - The overall hawkish/dovish stance - Substantive content or meaning ONLY EDIT: 1. **Vague temporal phrases**: sufficiently long duration -> Use context from same text to infer timeframe some time’, appropriate period -> Make more specific using surrounding discussion 2. **Internal contradictions**: If prepared statement and Q&A conflict -> Align using the clearer version 3. **Excessive hedging**: Multiple qualifiers (’may potentially consider’) -> Simplify if context supports IMPORTANT: Base any specificity ONLY on information already in the text. Do NOT invent dates, numbers, or commitments. ORIGINAL TEXT ({date}): {text} EDITED VERSION:
Note: Prompt used to generate counterfactual press conference transcripts with reduced linguistic ambiguity. The editor (ChatGPT-5o) is constrained to preserve all substantive policy content while only clarifying vague language, resolving contradictions, and reducing excessive hedging. Placeholders {date} and {text} are replaced with the actual conference date and transcript for each editing task.
This cross-LLM design is the pillar of our counterfactual analysis. By using different models for editing (ChatGPT 5o-mini) and evaluation (Gemini 2.5-Flash), we mitigate a critical concern: that the evaluator might mechanically align with the editor’s interpretation simply because both tasks are performed by the same model. In simple words, if we used a single LLM for both editing and evaluation, we would not be able to distinguish genuine clarity improvements from the model’s tendency to find its own edits more coherent, a form of confirmation bias at the algorithmic level to which LLMs are prone. The cross-model design ensures that disagreement reduction reflects genuine improvements in textual clarity rather than model-specific interpretive consistency.
6.6 Results
Figure 19 shows the ATEs. Panel (a) shows the mean disagreement change in basis points, while panel (b) displays the percentage change across the three OIS tenors.
Figure 19: Mean Disagreement Change by Tenor
(a) Basis Points Reduction
(b) Percentage Reduction
Note: Mean disagreement change following linguistic edits by ChatGPT, evaluated by Gemini 2.5-Flash across three OIS tenors (3-month, 2-year, and 10-year). Panel (a) shows the reduction in basis points, while panel (b) shows the percentage reduction in forecast disagreement (reduction = positive value). Error bars represent 95% confidence intervals from one-sided t-tests. Disagreement is measured as the cross-sectional standard deviation of rate forecasts across 30 simulated heterogeneous traders interpreting ECB press conference transcripts. Sample covers 30 highest-volatility ECB Governing Council meetings in the period 1998-2025.
The editing intervention achieves statistically significant reductions in disagreement at the 2Y and 10Y maturity, while the 3M tenor uncertainty slightly diminishes, but the impact is highly volatile across conferences. For the 10-year OIS, editing reduces disagreement by an average of 7.6 basis points (27.8% of the original volatility), while the 2-year tenor exhibits a reduction of 5.9 basis points (21.4%). The 3-month tenor shows a more modest, yet still meaningful, reduction of 2.1 basis points (4.44%). The maturity structure of these effects aligns with previous results. Longer-term rates are more sensitive to forward guidance and economic outlook assessments, which are precisely the textual elements most amenable to clarification. Short-term rates, by contrast, are more tightly anchored to the immediate policy decision and exhibit less scope for interpretive disagreement even in the baseline scenario.
Figure 20 visualizes the conference-by-conference pattern of disagreement change. The heatmap clearly identifies specific press conferences where editing achieved substantial reductions (shown in dark blue) versus those where the impact was minimal or even counterproductive (shown in dark red).
Figure 20: Change in Disagreement by Press Conference
Note: Percentage reduction in disagreement for each of the 30 conferences with the highest market-based disagreement, by tenor (3-month, 2-year, and 10-year). ChatGPT (5o-mini) edits the original ECB press conference transcripts to reduce linguistic ambiguity, then Gemini (2.5-Flash) agents evaluate both original and edited versions. Disagreement is measured as the cross-sectional standard deviation of rate forecasts across 30 heterogeneous agents. Darker colors indicate larger changes in disagreement. Negative values indicate cases where editing increased disagreement.
Two patterns emerge from the heatmap. First, the predominance of blue tiles indicates that editing reduces disagreement in the vast majority of conferences across all tenors. Second, when editing fails to reduce disagreement, or even increases it, these instances tend to occur consistently across all three maturities rather than being tenor-specific. The latter cases cluster in three distinct periods: the GFC (late 2008-early 2009), the recent tightening cycle (December 2022), and the early ECB period (late 1998), suggesting that LLM editing is less effective during episodes of extreme market stress or high institutional uncertainty. This result is consistent with a simple theoretical model of expectation formation under heterogeneous information processing. In the aforementioned setting, editing primarily reduces textual ambiguity by clarifying hedging language and resolving contradictions. Under normal conditions, this effectively compresses disagreement as agents converge toward a clearer public signal. However, during crisis episodes, agents enter the press conference with wildly divergent priors about the economic state and appropriate policy stance: when prior dispersion is so large, even clear communication cannot eradicate disagreement. The synchronous failures across all tenors confirm that this prior dispersion affects expectations uniformly across the yield curve, rather than reflecting tenor-specific uncertainty.
Several considerations inform the interpretation of these results. First, the editing protocol we employ is necessarily stylized. Real-world communication involves trade-offs between clarity and flexibility, and central banks may deliberately preserve ambiguity to avoid premature commitment or to accommodate internal disagreement among policymakers. Our simplified editing instructions abstract from these strategic considerations. Second, human evaluation rather than LLM based appraisal would constitute the gold standard for validating our findings. While our LLM-based measurement provides scalable and replicable disagreement estimates, we cannot definitively rule out that both the editor (ChatGPT) and evaluator (Gemini) share systematic biases in how they process certain linguistic features. A robust extension would involve having professional economists or central bank communication specialists assess: (i) whether the edited transcripts preserve the intended policy message, (ii) whether the edits are institutionally and legally permissible, and (iii) whether human readers perceive the edited versions as genuinely clearer. Such validation would confirm whether the disagreement reductions we observe reflect actual improvements in interpretability or merely artifacts of model-specific text processing. Until then, our results should be interpreted as proof-of-concept that linguistic editing can reduce measured disagreement, rather than as direct policy recommendations for communication strategy.